OpenStudy (yisrael012332):

Evaluate. cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5 5 to the power of negative 1 over 6 5 to the power of 3 over 2 5 to the power of 5 over 2 5 to the power of negative 5 over 6

OpenStudy (yisrael012332):

55√ −−−−−−√3 55 −−−−√3

OpenStudy (yisrael012332):

@jgirl128

OpenStudy (yisrael012332):

@plzzhelpme

OpenStudy (yisrael012332):

@BloomLocke367

OpenStudy (yisrael012332):

@BioHazard9064

OpenStudy (yisrael012332):

@DarkMoonZ

OpenStudy (yisrael012332):

@AaronAndyson

OpenStudy (yisrael012332):

@zzr0ck3r

OpenStudy (yisrael012332):

@UsukiDoll

OpenStudy (anonymous):

Is this the question?\[\frac{\sqrt[3]{5} \times \sqrt{5}}{\left( \sqrt[3]{5} \right)^5}\]

OpenStudy (mertsj):

\[\frac{5^{\frac{1}{3}}\times 5^{\frac{1}{2}}}{(5^{\frac{1}{3})^{5}}}=\frac{5^{\frac{5}{6}}}{5^{\frac{5}{3}}}=5^{-\frac{5}{6}}\]

OpenStudy (usukidoll):